A hospital staff member mixed a 65% disinfectant solution with a 15% disinfectant solution. How many liters of each were used to make 50 liters of a 37% disinfectant solution?
Answer by edjones(8007)

Let x=liters of 65% soln.
50-x= liters of 15% soln.
liters * % = amount of disinfectant
.65x+.15(50-x)=.37*50
.65x+7.5-.15x=18.5
subtract 7.5 from both sides.
.5x=11
divide .5x into both sides
x=22 liters of 65% soln.
50-x=28 liters of 15% soln.
Check:
(22*.65)+(28*.15)=18.5
EdJones